# CasualHDRSplat: Robust High Dynamic Range 3D Gaussian Splatting from Casually Captured Videos

### v1

@@ 3.3 Physical Image Formation Model
$\mathbf{B}(x,y)=F\left(\int_{t_{b}}^{t_{b}+\Delta t}\mathbf{H}\left(x,y,\mathbf{P}(t)\right)\mathrm{dt}\right)$ (7)
@@
$\mathbf{H}(x,y)\approx\frac{1}{N}\sum_{k=0}^{N-1}\mathbf{H}_{k}\left(x,y,\mathbf{P}(t_{k})\right)\Delta t$ (8)
@@
$\mathbf{B}(x,y)=\mathbf{F}(\mathbf{H}(x,y))=\text{TM}\circ\text{WB}(\mathbf{H}(x,y))$ (9)
$\text{WB}(\mathbf{c})=\begin{bmatrix}wb_{r},wb_{g},wb_{b}\end{bmatrix}^{T}\odot\begin{bmatrix}c_{r},c_{g},c_{b}\end{bmatrix}^{T}$ (10)
@@ 3.4 Loss Function
$\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda_{\text{exp}}\mathcal{L}_{\text{exp}},\quad\mathcal{L}_{\text{rec}}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}},$ (11)
where $\mathcal{L}_{\text{rec}}$ can constrain the consistency between the rendered image $\mathbf{C}_{k}(\mathbf{x})$ (the $k^{th}$ blurry LDR image synthesized from 3D-GS using the aforementioned image formation model (Eq. 8)) and the input LDR image $\mathbf{C}^{gt}_{k}(\mathbf{x})$.
To accurately model significant exposure variations in the input images, the second term of the loss function normalizes the images to a medium exposure by scaling pixel intensities before computing discrepancies [32, 53]:
$\small\mathcal{L}_{\text{exp}}=\mathcal{L}_{1}\left(\frac{\mathbf{C}^{gt}_{k}(\mathbf{x})}{\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})},\frac{\mathbf{C}_{k}(\mathbf{x})}{\bar{\mathbf{C}}_{k}(\mathbf{x})}\right)+\mathcal{L}_{\text{D-SSIM}}\left(\frac{\mathbf{C}^{gt}_{k}(\mathbf{x})}{\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})},\frac{\mathbf{C}_{k}(\mathbf{x})}{\bar{\mathbf{C}}_{k}(\mathbf{x})}\right),$ (12)
where $\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})$ and $\bar{\mathbf{C}}_{k}(\mathbf{x})$ represent the average pixel value of $\mathbf{C}^{gt}_{k}(\mathbf{x})$ and $\mathbf{C}_{k}(\mathbf{x})$. We set $\lambda_{\text{exp}}=0.25$ in all our experiments[20].
4 Experiments
Figure 3: Qualitative comparison on the Girls-vicon sequence of the Realsense dataset in terms of NVS.
Figure 4: Qualitative comparison on the Building sequence of the Smartphone dataset in terms of NVS.
@@ 4 Experiments
Figure 5: Qualitative comparison on the Trolley sequence of the synthetic dataset in terms of NVS.
Figure 6: Qualitative comparison on the Pool sequence of the synthetic dataset under training view.
4.1 Datasets
Synthetic datasets. We generated a synthetic dataset using Blender 3.6 with the Cycles engine, featuring four distinct scenes: Factory, Pool, Cozyroom, and Trolley. Each scene contains 77 images. The dataset generation combines physical motion blur imaging model[54] and tone mapping[13] from HDR to LDR. For each scene, images were assigned random exposure times, and captured using a continuous camera trajectory, generating sharp HDR images, which were averaged over the exposure period to create motion-blurred images.
Real datasets. Current related datasets use multi-view images with known exposure times from fixed viewpoints during exposure time, unlike our approach using casual videos for HDR reconstruction. Therefore, we collected a challenging real-world dataset, CasualVideo, using Intel RealSense D455 and Google Pixel 8 Pro on a DJI RS3 Mini gimbal. The dataset includes two subsets: RealSense and Smartphone. RealSense has four sequences (Yakitori, Toufu, Toufu-vicon, Girls-vicon), with ground truth poses for the latter two from Vicon. Smartphone has two sequences: Building and Fish. Since RealSense cannot provide current exposure times when auto-exposure is enabled, which limits baseline methods relying on exposure data, we implemented custom auto-exposure control with fixed aperture and ISO, following [48], on both cameras. To further evaluate our method on real-world data, we used the publicly available ScanNet[7], which features scenes recorded with auto-exposure enabled[2].
Table 1: Quantitative comparisons on the synthetic datasets in terms of novel view
@@ 4.1 Datasets
gsplat [18] 15.14 0.75 0.25 11.73 0.65 0.30 14.48 0.62 0.32 13.86 0.76 0.22
Gaussian-W [62] 23.68 0.75 0.26 23.28 0.69 0.62 17.83 0.64 0.34 27.16 0.85 0.15
BAD-Gaussians [64] 14.99 0.81 0.25 25.09 0.72 0.30 16.12 0.66 0.22 17.12 0.75 0.20
HDR-Plenoxels [17] 24.36 0.72 0.29 30.84 0.81 0.33 17.05 0.55 0.42 28.13 0.81 0.13
HDR-NeRF [13] 14.57 0.31 0.68 - - - - - - 13.62 0.32 0.77
CasualHDRSplat-random (ours) 30.25 0.89 0.10 32.63 0.91 0.09 25.14 0.81 0.24 29.62 0.86 0.10
CasualHDRSplat-gt (ours) 30.75 0.90 0.09 32.36 0.92 0.08 25.85 0.88 0.11 31.32 0.92 0.09
Table 2: Quantitative comparisons on the real-world datasets in terms of novel view.
@@
CasualHDRSplat-random (ours) 28.30 0.83 0.13 28.79 0.83 0.09 30.87 0.90 0.15 32.00 0.90 0.19
CasualHDRSplat-gt (ours) 30.81 0.87 0.12 29.71 0.85 0.08 31.34 0.92 0.12 32.39 0.91 0.17
4.2 Implementation Details
@@
We implemented our method in PyTorch using the gsplat[59] with MCMC strategy[19]. The optimization of all the parameters was done using the Adam optimizer, keeping the Gaussian primitive learning rate consistent with gsplat. To balance performance and efficiency, we set the number of virtual camera poses (n in Eq. 8) to 10.
@@
For initialization, we used HLoc[46] on our Sythetic and Realsense dataset .For the ScanNet[7] and our Smartphone datasets, we used DPV-SLAM [31] since HLoc failed due to poor image quality. We use NVIDIA RTX 3090 on synthetic data and A6000 on real data, respectively.
4.3 Baseline Methods and evaluation metrics
@@
To evaluate the robustness of our method in learning accurate scenes representation under poorly exposed conditions and server motion blur, we compared it against scene reconstruction methods that handle brightness variations, e.g. HDR-NeRF[13], HDR-Plenoxels[17], Gaussian-W[62], as well as method for scene reconstruction from blurred images, such as BAD-Gaussians[64]. In addition, 3D-GS[18] implemented by gsplat[59] was included as the comparison baseline. Following previous works[17, 4, 13], we evaluate LDR images rendered from the learned scene used metrics such as PSNR, SSIM[57], and LPIPS[63]. Furthermore, to evaluate whether our method effectively recovers camera motion trajectories, we compared it with pose estimation method, e.g. HLoc[46], DPV-SLAM[31], BAD-Gaussians[64]. For pose estimation accuracy, we utilize absolute trajectory error (ATE) with mean and std as the metric. Since $I^{2}$-SLAM[2] and HDR-GS[4] are not open-sourced, we cannot compare our method against them.
4.4 Quantitative evaluation results.
@@
We conducted experiments with two settings: randomly initialized exposure times (CasualHDRSplat-random) and ground truth exposure times (CasualHDRSplat-gt). We evaluated our method’s performance on novel view synthesis, image deblurring, and pose estimation (ATE) tasks. More results on the ScanNet dataset and two Realsense scenes can be found in the supplementary materials.
@@
Due to the blurry nature of most real-world images, we used 5 to 10 sharp images per sequence for evaluation. As shown in Table 1 and Table 2, our method significantly outperforms previous methods in novel view synthesis. CasualHDRSplat-random, even with randomly initialized exposure times, surpasses prior works by jointly optimizing exposure times and CRF representation. Unlike HDR-NeRF[13], our method reconstructs HDR scenes accurately from degraded images without measured exposure times, whereas HDR-NeRF fails on all real dataset scenes.
@@
By modeling the physical principles of camera imaging, our method also outperforms HDR-Plenoxels[17] and Gaussian-W[62]. Moreover, our approach leverages spline representations for optimizing camera motion trajectories, enabling effective scene learning, which other methods struggle with in the absence of ground truth poses.
@@
Table 3: Ablation studies on the $ratio$ for initializing camera motion spline.
@@
Table 4: Ablation study: Average results across two datasets to investigate the effect on model performance.
@@
Table 5: Quantitative comparisons on the synthetic datasets in terms of deblur
Factory Pool Trolley Cozyroom
[…]
BAD-Gaussians[64] 24.32 0.73 0.12 25.87 0.79 0.23 19.06 0.62 0.19 23.37 0.79 0.11
CasualHDRSplat-random (ours) 31.20 0.88 0.05 32.95 0.87 0.10 23.65 0.69 0.12 29.60 0.84 0.05
CasualHDRSplat-gt (ours) 32.00 0.91 0.07 34.53 0.96 0.05 29.35 0.87 0.08 33.01 0.93 0.04
Table 6: Quantitative comparisons for pose estimation on the Realsense sequences with Vicon motion captured groundtruth. The results are in the absolute trajectory error metric (ATE) with units in centimeters.
HLoc DPV-SLAM BAD-GS CasualHDRSplat-random CasualHDRSplat-gt
Toufu-vicon .4644$\pm$.3921 .4043$\pm$.3877 .3935$\pm$.4212 .3687$\pm$.3874 .3595$\pm$.3462
Girls-vicon $1$9557$\pm$.8231 .8548$\pm$.8628 .8294$\pm$.8834 .6478$\pm$.8268
@@
Table 5 shows that our method achieves superior performance in image deblurring task compared to BAD-Gaussians[64]. This is because our method can recover accurate scene representation from images affected by both motion blur and poor exposure.
@@
The experimental results presented in Table 6 demonstrate that our method outperforms prior approaches in the pose estimation task. HLoc, which relies on feature point matching, exhibits poor performance under conditions of varying brightness and motion blur. Although DPV-SLAM[31] and BAD-Gaussians[64] can operate effectively in the presence of motion blur, they struggle to tolerate environments with high-contrast and varying exposure time. This indicates that our method can robustly estimate continuous camera trajectories under high-contrast environments, within varying exposure time and motion blur.
4.5 Qualitative evaluation results.
@@
The qualitative comparisons of the NVS and deblurring tasks on both synthetic and real datasets are shown in Figure 3, 4, 5 and 6. The experimental results indicate that our method outperforms previous approaches. This demonstrates that our method can effectively learn scene representations from images that simultaneously exhibit varying exposure time and motion blur, while prior works lack robustness given the challenging conditions and failed to reconstruct high-quality 3D HDR scene. Note that BAD-Gaussians can deblur the training views as ours, however, due to its failure of pose optimization, the rendered images are misaligned with others.
4.6 Ablation studies.
@@
Initialization for camera motion spline. In our method, the camera motion spline needs to be initialized by leveraging the poses estimated from HLoc [46] or DPV-SLAM [31]. Therefore, the configuration of initialization will impact the performance of our method. We define a ratio representing the number of control knots of spline divided by the number of input images and evaluate the effect of the ratio. The results in Table 4 indicate that model performance improves until it saturates as the ratio increases. We set $\textit{ratio}=3.0$ for all experiments to ensure a trade-off between the performance and computational overhead.
@@
5 Conclusion
In this paper, we introduce a novel method CasualHDRSplat for reconstructing 3D HDR scenes from casually captured videos, which often exhibit limited dynamic range and motion blur. Our method reconstruct 3D HDR scene and generate LDR images given specified exposures and camera poses, providing high robustness and flexibility. By leveraging the camera auto-exposure, we incorporate the latent high dynamic range of captured videos into a unified physical imaging model. This allows for the joint optimization of exposure time, continous-time camera trajectory, and CRF, enabling accurate 3D HDR scene reconstruction. Extensive experiments demonstrate that our method outperforms previous approaches in 3D HDR reconstruction.
[…]
Supplementary Material
6 Appendix

### v2

@@ 3.3. Physical Image Formation Model
(7) $\mathbf{B}(x,y)=F\left(\int_{t_{b}}^{t_{b}+\Delta t}\mathbf{H}\left(x,y,\mathbf{P}(t)\right)\mathrm{dt}\right)$
@@
(8) $\mathbf{H}(x,y)\approx\frac{1}{N}\sum_{k=0}^{N-1}\mathbf{H}_{k}\left(x,y,\mathbf{P}(t_{k})\right)\Delta t$
@@
(9) $\mathbf{B}(x,y)=\mathbf{F}(\mathbf{H}(x,y))=\text{TM}\circ\text{WB}(\mathbf{H}(x,y))$
(10) $\text{WB}(\mathbf{c})=\begin{bmatrix}wb_{r},wb_{g},wb_{b}\end{bmatrix}^{T}\odot\begin{bmatrix}c_{r},c_{g},c_{b}\end{bmatrix}^{T}$
@@ 3.4. Loss Function
(11) $\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda_{\text{exp}}\mathcal{L}_{\text{exp}},\quad\mathcal{L}_{\text{rec}}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}},$
where $\mathcal{L}_{\text{rec}}$ can constrain the consistency between the rendered image $\mathbf{C}_{k}(\mathbf{x})$ (the $k^{th}$ blurry LDR image synthesized from 3D-GS using Eq. 8) and the input LDR image $\mathbf{C}^{gt}_{k}(\mathbf{x})$.
To accurately model significant exposure variations in the input images, the second term of the loss function normalizes the images to a medium exposure by scaling pixel intensities before computing discrepancies (Liu et al., 2020; Wang et al., 2024b):
(12) $\small\mathcal{L}_{\text{exp}}=\mathcal{L}_{1}\left(\frac{\mathbf{C}^{gt}_{k}(\mathbf{x})}{\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})},\frac{\mathbf{C}_{k}(\mathbf{x})}{\bar{\mathbf{C}}_{k}(\mathbf{x})}\right)+\mathcal{L}_{\text{D-SSIM}}\left(\frac{\mathbf{C}^{gt}_{k}(\mathbf{x})}{\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})},\frac{\mathbf{C}_{k}(\mathbf{x})}{\bar{\mathbf{C}}_{k}(\mathbf{x})}\right),$
where $\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})$ and $\bar{\mathbf{C}}_{k}(\mathbf{x})$ represent the average pixel value of $\mathbf{C}^{gt}_{k}(\mathbf{x})$ and $\mathbf{C}_{k}(\mathbf{x})$. We set $\lambda_{\text{exp}}=0.25$ in all our experiments.
4. Experiments
@@
Figure 5. Qualitative comparison on the Trolley sequence of the synthetic dataset in terms of NVS.
Figure 6. HDR editing with various designated exposure times.
4.1. Datasets
Synthetic datasets. We created a synthetic dataset in Blender 3.6 (Cycles engine) with four photorealistic scenes: Factory, Pool, Cozyroom, and Trolley, each containing 77 images with realistic lighting and geometry. Following Wang et al. (2023); Huang et al. (2022), we combined a physical imaging model and tone-mapping. Images were rendered along continuous camera trajectories with random exposure times, producing sharp HDR frames that were temporally averaged to simulate motion-blurred LDR frames.
Real datasets. Existing HDR datasets use static-view stacks with known exposures, differing from our casual video-based setting. To bridge this gap, we collected a real-world dataset, CasualVideo, using Intel RealSense D455 and Google Pixel 8 Pro on a DJI RS3 Mini gimbal. It includes two subsets: RealSense (4 sequences, 2 with Vicon poses) and Smartphone (2 sequences). As RealSense lacks exposure metadata under AE, we implemented custom AE control (fixed aperture/ISO) on both devices, following Su and Kuo (2015). We also evaluate on ScanNet (Dai et al., 2017), which uses auto-exposure (AE)  (Bae et al., 2024).
Table 1. Quantitative comparisons on the synthetic datasets in terms of novel view
@@ 4.1. Datasets
gsplat (Kerbl et al., 2023) 15.14 0.75 0.25 11.73 0.65 0.30 14.48 0.62 0.32 13.86 0.76 0.22
Gaussian-W (Zhang et al., 2024) 23.68 0.75 0.26 23.28 0.69 0.62 17.83 0.64 0.34 27.16 0.85 0.15
BAD-Gaussians (Zhao et al., 2024) 14.99 0.81 0.25 25.09 0.72 0.30 16.12 0.66 0.22 17.12 0.75 0.20
HDR-Plenoxels (Jun-Seong et al., 2022) 24.36 0.72 0.29 30.84 0.81 0.33 17.05 0.55 0.42 28.13 0.81 0.13
HDR-NeRF (Huang et al., 2022) 14.57 0.31 0.68 - - - - - - 13.62 0.32 0.77
Casual3DHDR-random (ours) 30.25 0.89 0.10 32.63 0.91 0.09 25.14 0.81 0.24 29.62 0.86 0.10
Casual3DHDR-gt (ours) 30.75 0.90 0.09 32.36 0.92 0.08 25.85 0.88 0.11 31.32 0.92 0.09
Table 2. Quantitative comparisons on the real-world datasets in terms of novel view.
@@
Casual3DHDR-random (ours) 28.30 0.83 0.13 28.79 0.83 0.09 30.87 0.90 0.15 32.00 0.90 0.19
Casual3DHDR-gt (ours) 30.81 0.87 0.12 29.71 0.85 0.08 31.34 0.92 0.12 32.39 0.91 0.17
4.2. Implementation Details
@@
We implemented our method in PyTorch using the gsplat (Ye et al., 2024) with MCMC strategy (Kheradmand et al., 2024). The optimization of all parameters was done using the Adam optimizer, keeping the Gaussian primitive learning rate consistent with gsplat. To balance performance and efficiency, we set the number of virtual camera poses (n in Eq. 8) to 10.
@@
For initialization, we used HLoc (Sarlin et al., 2019) on our Sythetic and Realsense dataset. For the ScanNet (Dai et al., 2017) and our Smartphone datasets, we used DPV-SLAM (Lipson et al., 2024) since HLoc failed due to poor image quality. Synthetic dataset experiments were conducted on an NVIDIA RTX 3090, while real dataset experiments used an NVIDIA RTX A6000 due to higher memory requirements.
4.3. Baseline Methods and evaluation metrics
@@
To evaluate the robustness of our method in learning accurate scenes representation under poorly exposed conditions and server motion blur, we compared it against scene reconstruction methods that handle brightness variations, e.g. HDR-NeRF (Huang et al., 2022), HDR-Plenoxels (Jun-Seong et al., 2022), Gaussian-W (Zhang et al., 2024), as well as method for scene reconstruction from blurred images, such as BAD-Gaussians (Zhao et al., 2024). In addition, 3D-GS (Kerbl et al., 2023) implemented by gsplat (Ye et al., 2024) was included as the comparison baseline. Following previous works (Jun-Seong et al., 2022; Cai et al., 2024; Huang et al., 2022), we evaluate LDR images rendered from the learned scene used metrics such as PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018). Furthermore, to evaluate whether our method effectively recovers camera motion trajectories, we compared it with pose estimation method, e.g. HLoc (Sarlin et al., 2019), DPV-SLAM (Lipson et al., 2024), BAD-Gaussians (Zhao et al., 2024). For pose estimation, we utilize absolute trajectory error (ATE) with mean and std as the metric.
4.4. Quantitative evaluation results.
@@
We conducted experiments with two settings: one using randomly initialized exposure times (Casual3DHDR-random) and the other with ground truth exposure times (Casual3DHDR-gt). We evaluated our method’s performance on novel view synthesis, image deblurring, and pose estimation tasks. More results on the ScanNet and Realsense scenes can be found in the supplementary material.
@@
Due to the blurry nature of most real-world images, we used 5 to 10 sharp images per sequence for evaluation. As shown in Table 1 and Table 2, our method significantly outperforms previous methods in novel view synthesis. Casual3DHDR-random, even with randomly initialized exposure times, surpasses prior works by jointly optimizing exposure times and CRF representation. Unlike HDR-NeRF (Huang et al., 2022), our method reconstructs HDR scenes accurately from degraded images without measured exposure times, whereas HDR-NeRF fails on all real dataset scenes.
@@
By modeling physical principles of imaging, our method also outperforms HDR-Plenoxels (Jun-Seong et al., 2022) and Gaussian-W (Zhang et al., 2024). Moreover, our approach leverages spline representations for optimizing camera motion trajectories, enabling effective scene learning, which other methods struggle with the absence of ground truth poses.
Table 3. Quantitative comparisons on the synthetic datasets in terms of deblur
Factory Pool Trolley Cozyroom
[…]
BAD-Gaussians (Zhao et al., 2024) 24.32 0.73 0.12 25.87 0.79 0.23 19.06 0.62 0.19 23.37 0.79 0.11
Casual3DHDR-random (ours) 31.20 0.88 0.05 32.95 0.87 0.10 23.65 0.69 0.12 29.60 0.84 0.05
Casual3DHDR-gt (ours) 32.00 0.91 0.07 34.53 0.96 0.05 29.35 0.87 0.08 33.01 0.93 0.04
Table 4. Quantitative comparisons on the ScanNet dataset in term of deblur (using BRISQUE metric).
BAD-Gaussians (Zhao et al., 2024) 42.78 60.48 57.37 42.72 67.30 65.01 55.94
Casual3DHDR-random (ours) 38.08 47.65 53.37 37.55 62.35 58.23 49.53
Table 5. Quantitative comparisons for pose estimation on the Realsense sequences with Vicon motion captured groundtruth. The results are in the absolute trajectory error metric (ATE) with units in centimeters.
HLoc (Sarlin et al., 2019) DPV-SLAM (Lipson et al., 2024) BAD-Gaussians (Zhao et al., 2024) Casual3DHDR-random Casual3DHDR-gt
[…]
@@
Table 6. Ablation studies on the interpolation ratioratio for initializing camera motion spline.
@@
Table 7. Ablation study: Average results across two datasets to investigate the effect on model performance.
@@
Figure 7. Qualitative comparison on the Pool sequence of the synthetic dataset under training view. BAD-Gaussians is capable to deblur the training views as ours Due to the failure of pose optimization in the BAD-Gaussians, its image are misaligned with others.
@@
Table 3 shows that our method achieves superior performance in image deblurring task compared to BAD-Gaussians (Zhao et al., 2024). This is because our method can recover accurate scene representation from images affected by both motion blur and poor exposure.
In addition, we also evaluated the quantitative metrics for deblurring on public real-world datasets (i.e. ScanNet (Dai et al., 2017)). Due to most images of ScanNet dataset are motion-blurred, we can not find sharp reference images for evaluating, thus we utilize the no-reference image quality metric BRISQUE (Mittal et al., 2012) to quantitatively compare the deblurring performance between our method and BAD-Gaussians (Zhao et al., 2024), as shown in Table 4.
@@
The experimental results presented in Table 5 demonstrate that our method outperforms prior approaches in the pose estimation task. HLoc, which relies on feature point matching, exhibits poor performance under conditions of varying brightness and motion blur. Although DPV-SLAM (Lipson et al., 2024) and BAD-Gaussians (Zhao et al., 2024) can operate effectively in the presence of motion blur, they struggle to tolerate environments with high-contrast and varying exposure time. This indicates that our method can robustly estimate continuous camera trajectories under high-contrast environments, within varying exposure time and motion blur.
4.5. Qualitative evaluation results.
@@
The results in Figure 6 demonstrate that our method can accurately recover HDR scenes, while the brightness of the rendered images can be manually adjusted by changing the exposure time. Qualitative comparisons of the NVS and deblurring tasks on both synthetic and real datasets are shown in Figure 3, Figure 4, Figure 5 and Figure 7. The experimental results indicate that our method outperforms previous approaches and is visually closer to the ground truth. This demonstrates that our method can effectively recover high-quality 3D HDR scene from images that simultaneously exhibit varying exposure time and motion blur, while prior work lacks robustness given the challenging conditions.
4.6. Ablation studies.
@@
Initialization for camera motion spline. In our method, the camera motion spline needs to be initialized by leveraging the poses estimated from HLoc (Sarlin et al., 2019) or DPV-SLAM (Lipson et al., 2024). Therefore, the configuration of initialization will impact the performance of our method. We define a ratio of interpolation representing the number of control knots of spline divided by the number of input images and evaluate the effect of the ratio. The results in Table 7 indicate that model performance improves until it saturates as the ratio increases. We set $\textit{ratio}=3.0$ for all experiments to ensure a trade-off between the performance and computational overhead.
@@
5. Conclusion
In this paper, we present Casual3DHDR, a novel framework for reconstructing high-quality 3D HDR scenes from casually-captured auto-exposure (AE) videos, which often exhibit limited dynamic range, motion blur, and uncontrolled camera motion. Our approach effectively reconstructs 3D HDR scenes and generates photorealistic LDR images for specified exposures and camera poses, demonstrating robustness and versatility. By integrating latent motion and temporal dynamic range from AE videos into a unified physical imaging model, our method jointly optimizes exposure time, continuous camera trajectories, and the camera response function (CRF). Extensive experiments validate the accuracy and effectiveness of Casual3DHDR, highlighting its potential to advance 3D HDR scene reconstruction from challenging real-world inputs.
6. Acknowldgement
This work was supported in part by the National Natural Science Foundation of China (Grants 62202389, 42371451, 42394061), in part by the Natural Science Foundation of Wuhan (No.2024040701010028), in part by a grant from the Westlake University-Muyuan Joint Research Institute and in part by the Westlake Education Foundation. We specially thank Xiang Liu for his kind and valuable suggestions during the writing of this paper.
Appendix A Appendix